# Spectraformer: A Unified Random Feature Framework for Transformer

### v2

@@ 2.3 Random Features for Kernel Approximation
Of the kernel approximation family of methods, the most successful and appropriate for linearized attention in Transformer has been the random features approach [19]. Estimating kernels using random features relies on a fundamental insight from harmonic analysis called Bochner’s theorem [20] stated as follows:
A continuous shift-invariant kernel $\mathcal{K}(\boldsymbol{x},\boldsymbol{y})=\mathcal{K}(\boldsymbol{x}-\boldsymbol{y})$ on $\mathbb{R}^{d}$ is positive definite if and only if $\mathcal{K}(\boldsymbol{\delta})$ is the Fourier transform of a non-negative measure $p(\boldsymbol{\omega})$. If $\mathcal{K}(\boldsymbol{\delta})$ is properly scaled, that is $\mathcal{K}(0)=1$ then the measure $p(\boldsymbol{\omega})$ is a proper probability distribution.
@@
$\mathcal{K}(\boldsymbol{\delta})=\int p(\boldsymbol{\omega})\exp(i\boldsymbol{\omega\delta})\mathrm{d}\boldsymbol{\omega}$ (9)
When approximating the integration in this equation via Monte Carlo sampling with $s$ samples, $\boldsymbol{\omega}\sim p(\boldsymbol{\omega})$, we can obtain the feature map:
$\begin{array}[]{ll}\mathcal{K}(\boldsymbol{\delta})&=\mathbb{E}[\phi_{\boldsymbol{\omega}}(\boldsymbol{x})^{T}\phi_{\boldsymbol{\omega}}(\boldsymbol{y})]\approx<\phi_{\boldsymbol{\omega}}(\boldsymbol{x}),\phi_{\boldsymbol{\omega}}(\boldsymbol{y})>\\ \phi_{\boldsymbol{\omega}}(\boldsymbol{x})\!&=\!\frac{1}{\sqrt{s}}\big{[}\exp(-\mbox{i}{\boldsymbol{\omega}}^{\!\top}_{1}\boldsymbol{x}),\ldots,\exp(-\mbox{i}{\boldsymbol{\omega}}^{\!\top}_{s}\boldsymbol{x})]^{\!\top}\end{array}$ (10)
@@ 2.4 Random Features for Kernel Learning
Previous random feature techniques are robust in approximating kernels for specific learning problems. However, kernels may be chosen using heuristics and by convention from a popular subset. Kernel is very much a hyperparameter. This could pose a problem since there is no free lunch in learning as much as in kernel choice [21]. [4] has shown experimental results that this is also true in the context of attention in Transformer. Hence, instead of picking a kernel for the attention, we could learn the kernel via kernel learning an established kernel methods technique. Kernel learning has shown to be effective in general [22, 23, 24] and in the case of attention via Gaussian Mixture Model (GMM), learnable weight matrices (FastFood), deep generative model (DGM) [14]. Kernel learning is achieved by making the weight matrix $\boldsymbol{W}$ learnable where $\boldsymbol{\omega}_{i}$ is the output of a function parameterizing $p(\boldsymbol{\omega})$.
@@ 3 Spectraformer
Spectraformer is a unifed random feature framework that allows for the combination of any weight matrix with any component function to be used as kernelized attention in Transformer. The general equation for random feature is shown in Equation 11, generalized from Equation 10, based on Liu et al. [7], Choromanski et al. [8].
$\begin{array}[]{rl}\phi_{\boldsymbol{\omega}}(\boldsymbol{x})=&\frac{1}{\sqrt{s}}\big{[}a_{1}f_{1}(\boldsymbol{\omega}_{1},\boldsymbol{x}),...,a_{s}f_{1}(\boldsymbol{\omega}_{s},\boldsymbol{x}),...,\\ &a_{1}f_{l}(\boldsymbol{\omega}_{1},\boldsymbol{x}),...,a_{s}f_{l}(\boldsymbol{\omega}_{s},\boldsymbol{x})]^{\!\top}\\ \boldsymbol{W}=&[\boldsymbol{\omega}_{1},...,\boldsymbol{\omega}_{s}]^{T}\in\mathbb{R}^{s*d}\end{array}$ (11)
Spectraformer is shown in Figure 2. The attention of the OT (on the left-hand side) is replaced by the attention of the Spectraformer which pre-computes the Hadamard (element-wise) product of values $\boldsymbol{v}$ and transformed keys $\phi(\boldsymbol{k})$ scaled by $\phi(\boldsymbol{k})$, the product of which is multiplied with the transformed queries $\phi(\boldsymbol{q})$. The linear map $\phi$ is composed from weight matrices and component functions. The linear map has $s$ components with $a=1$ (see Equation 11).
The figure and the equation together highlight the three strands of research as mentioned previously in Section 1: (A) Weight matrix approximator: constructing $\boldsymbol{W}$ more effectively; (B) Weight matrix learner: parameterizing $\boldsymbol{W}$ instead of sampling; (C) Component function: constructing $f_{j}$ more effectively.
@@ 3.1 Weight Matrices
In Spectraformer, the weight matrix $\boldsymbol{\boldsymbol{W}}$ which is either an approximator (from the families of (e.g., ORF [9], SORF [9]), Unit-cube (e.g., QMC [25], MM [26]), Quadrature (e.g., SGQ [27]). or a learner (FastFoodL [14]).
@@
We term the matrix $\boldsymbol{W}$ as weight matrix instead of random matrix since although $\boldsymbol{W}$ acts like a random matrix ($\boldsymbol{\omega}_{i}\sim p(.)$, see Equation 10), it is not guaranteed to be one. In the case of weight matrix approximator, they approximate a random matrix and in the case of weight matrix learner, they parameterize a weight matrix, thereby implicitly learning a distribution $p(.)$ associated with such $\boldsymbol{W}$. The weight matrix is given in Definition 3.1.
@@
Provided that a matrix can substitute for a random matrix in Equation 11, and the solution for this equation given such matrix and $f$ as TrigRF being the solution for Equation 9, i.e., as an unbiased estimator of $\mathcal{K}$, and TrigRF being $f_{1}=f_{l}=\exp(-i\boldsymbol{\omega}^{T}\boldsymbol{x})$ (see Equation 10) then such a matrix is a weight matrix.
@@
It follows from Definition 3.1 that for any weight matrix $\boldsymbol{W}$, $f=TrigRF$ is a valid component function. Spectraformer enables weight matrix approximation in terms of three families based on how they solve the intractable integral in Equation 9:
@@
  • Monte Carlo sampling-based weight matrix involves approximating the integral of Equation 9 using Monte Carlo sampling with the solution given by Equation 10 [19] with $\boldsymbol{W}$ being the ‘Base’ random matrix of $p(.)$, $\boldsymbol{\omega}_{i}\sim p(.)$. The random matrix $\boldsymbol{W}$ can be constructed more efficiently (either reducing time or space complexity) with FastFood [28]. The random matrix $\boldsymbol{W}$ can also be enforced to form geometrical couplings such as orthogonality to reduce approximation error with ORF, SORF [9]. Both these methods produce valid random matrices.
@@
  • Unit-cube sampling-based weight matrix transforms Equation 9 to an integral on the unit cube $[0,1]^{d}$, then performs the approximation with uniform and independent points, thus reducing variance. The first approach is QMC [25], where $\boldsymbol{W}:\boldsymbol{\omega}_{i}=\Phi^{-1}(\boldsymbol{t}_{i})$, $\Phi$ being the cumulative distribution function (CDF) associated with $p(.)$ and $\boldsymbol{t}_{i}$ being a low discrepancy sequence. $\Phi^{-1}(\boldsymbol{t}_{i})\sim p(.)$ whilst having a lower variance than direct sampling from $p(.)$. MM [26, 7] improves over QMC by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$ on $\Phi^{-1}$.
@@
  • Quadrature-based weight matrix uses quadrature rules with non-uniform deterministic weights. The main approach explored is SGQ [27] which uses one-dimensional Gaussian quadrature rule by assuming $\mathcal{K}$ factorizes with respect to the dimensions. Smolyak rule is further implemented to alleviate the curse of dimensionality.
@@
[7] show that in various kernel benchmark datasets (among combination with TrigRF), QMC and ORF have consistently high performance. Due to the significant number of weight matrices in random features literature, we decide to cover only the most prominent ones. Appendix A.3 provides the technical details of the approximators discussed here and those that we left out.
@@
Spectraformer also allows to learn weight matrices by learning $p(.)$ via parameterizing $\boldsymbol{W}$. [29] introduce several of these parametrization schemes, which are then adapted into the attention setting by [14]. The best performing method is via making the constituent matrices of FastFood learnable, which we term FastFoodL. In general, a weight matrix learner is valid as long as it can implicitly model any distribution. In addition the TrigRF, PosRF has also been applied on FastFoodL. Technical details and other learners we have not covered are explored in Appendix A.4. We now explore existing works which have developed from Rahimi and Recht [19] (see Equation 9, 10) and how they lead to the development of Spectraformer.
@@ 3.2 Component functions
Spectraformer incorporates component functions $f$ such as PosRF, OPRF, SADERF. Component functions in Spectraformer combine each weight matrix row $\boldsymbol{\omega}_{i}$ with the input $\boldsymbol{x}$. The base case of the component function is TrigRF. Component function is given in Definition 3.2
@@
A function of $\boldsymbol{x}$ with respect to $\boldsymbol{\omega}$ is a valid component function if and only if $\mathcal{K}\approx\phi_{\boldsymbol{\omega}}(\boldsymbol{x})^{T}\phi_{\boldsymbol{\omega}}(\boldsymbol{y})$ given that $\boldsymbol{\omega}_{i}\sim p(.)$ in Equation 10.
@@
It follows that all component functions have provided theoretical guarantees given that $\boldsymbol{W}$ is a random matrix. In addition, PosRF, OPRF, and SADERF have also been proven theoretically to approximate $\mathcal{K}_{RBF}$ given $\boldsymbol{W}$ being the ORF.
@@
While TrigRF[19] is the base component function and guarantees $\mathcal{K}(\boldsymbol{\delta})$ to be close to $\phi_{TrigRF}(\boldsymbol{x})\cdot\phi_{TrigRF}(\boldsymbol{y})$ due to uniform convergence, it is typically evaluated using only the real component of $\exp(-i\boldsymbol{\omega}_{j}^{T}\boldsymbol{x})$. This leads to $\cos$ and $\sin$ functions which can produce unstable behavior, especially when $K(\boldsymbol{\delta})$ is close to 0, where variance is infinite [8, 12]. Therefore, TrigRF is avoided in Spectraformer.
@@
  • PosRF [8] is introduced in Spectraformer as a simple modification over TrigRF to counter this instability. PosRF guarantees positive function output and its variance approaching 0 as $K(\boldsymbol{\delta})$ approaches 0. [12] generalize TrigRF and PosRF to GERF.
@@
  • OPRF [12] is one of the solution to the reduction of the variance equation of GERF. OPRF is both positive and bounded (whereas PosRF is positive and unbounded). [13] generalize GERF to dense matrix instead of scalar parameters leading to DERF.
@@
  • SADERF [13] is one of the solutions to the reduction of the variance equation of DERF. Whilst DERF has many solutions, they all rely on operations which are incompatible with many deep learning libraries on GPU and TPU like SVD and eigen decompositions. SADERF is an extension of GERF with a tighter variance bound on DERF.
@@
The alternatives for component functions allow Spectraformer to approximate the RBF with $p(.)$ corresponding to the Gaussian. [13] and [12] show that SADERF-ORF is the SOTA combination followed by OPRF-ORF and PosRF-ORF. Technical details and other component function(s) not covered are explored in Appendix A.5.
@@
Table 1: Summary of research gap in kernelized attention. Rows are the component functions, columns are the weight matrices. For each cell, the left item is the cited work for this combination for typical kernel tasks, while the right item is the cited work for Transformer-based tasks. $\times$ indicates no prior work for the combination. Spectraformer explores all combinations listed here, excluding ‘Base’ due to exhaustive work as seen in the table.
f Monte Carlo Unit-cube Quadrature Learner
Base ORF SORF QMC MM SGQ FastFoodL
TrigRF [19], [8] [9], [8] [9], $\times$ [25], $\times$ [26], $\times$ [27], $\times$ [29], [14]
PosRF [8], [8] [8], [8] $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, [14]
OPRF [12], [12] [12], [12] $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$
SADERF [13], [13] [13], [13] $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$ $\times$, $\times$

### v3

@@ 2.3 Random Features for Kernel Approximation
Of the kernel approximation family of methods, the most successful and appropriate for linearized attention in Transformer has been the random features approach (Rahimi & Recht, 2007). Estimating kernels using random features relies on a fundamental insight from harmonic analysis called Bochner’s theorem (Rudin, 1990) stated as follows:
A continuous shift-invariant kernel $\mathcal{K}(\bm{x},\bm{y})=\mathcal{K}(\bm{x}-\bm{y})$ on $\mathbb{R}^{d}$ is positive definite if and only if $\mathcal{K}(\bm{\delta})$ is the Fourier transform of a non-negative measure $p(\bm{\omega})$. If $\mathcal{K}(\bm{\delta})$ is properly scaled, that is $\mathcal{K}(0)=1$ then the measure $p(\bm{\omega})$ is a proper probability distribution.
@@
$\mathcal{K}(\bm{\delta})=\int p(\bm{\omega})\exp(i\bm{\omega\delta})\mathrm{d}\bm{\omega}$ (9)
When approximating the integration in this equation via Monte Carlo sampling with $s$ samples, $\bm{\omega}\sim p(\bm{\omega})$, we can obtain the feature map:
$\begin{array}[]{ll}\mathcal{K}(\bm{\delta})&=\mathbb{E}[\phi_{\bm{\omega}}(\bm{x})^{T}\phi_{\bm{\omega}}(\bm{y})]\approx<\phi_{\bm{\omega}}(\bm{x}),\phi_{\bm{\omega}}(\bm{y})>\\ \phi_{\bm{\omega}}(\bm{x})\!&=\!\frac{1}{\sqrt{s}}\big{[}\exp(-\mbox{i}{\bm{\omega}}^{\!\top}_{1}\bm{x}),\ldots,\exp(-\mbox{i}{\bm{\omega}}^{\!\top}_{s}\bm{x})]^{\!\top}\end{array}$ (10)
@@ 2.4 Random Features for Kernel Learning
Previous random feature techniques are robust in approximating kernels for specific learning problems. However, kernels may be chosen using heuristics and by convention from a popular subset. The kernel is very much a hyperparameter. This could pose a problem since there is no free lunch in learning as much as in kernel choice (Schölkopf & Smola, 2018). Tsai et al. (2019) has shown experimental results that this is also true in the context of attention in Transformer. Hence, instead of picking a kernel for the attention, we could learn the kernel via kernel learning an established kernel methods technique. Kernel learning has shown to be effective in general (Tompkins et al., 2019; Wilson & Adams, 2013; Oliva et al., 2016) and in the case of attention via Gaussian Mixture Model (GMM), learnable weight matrices (FastFood), deep generative model (DGM) (Chowdhury et al., 2022). Kernel learning is achieved by making the weight matrix $\bm{W}$ learnable where $\bm{\omega}_{i}$ is the output of a function parameterizing $p(\bm{\omega})$.
@@ 3 Spectraformer
Spectraformer is a unified random feature framework that allows for the combination of any weight matrix with any component function to be used as kernelized attention in Transformer. The general equation for random feature is shown in Equation 11, generalized from Equation 10, based on Liu et al. (2022); Choromanski et al. (2021).
$\begin{array}[]{rl}\phi_{\bm{\omega}}(\bm{x})=&\frac{1}{\sqrt{s}}\big{[}a_{1}f_{1}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{1}(\bm{\omega}_{s},\bm{x}),...,\\ &a_{1}f_{l}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{l}(\bm{\omega}_{s},\bm{x})]^{\!\top}\\ \bm{W}=&[\bm{\omega}_{1},...,\bm{\omega}_{s}]^{T}\in\mathbb{R}^{s*d}\end{array}$ (11)
Spectraformer is shown in Figure 1. The attention of the OT (on the left-hand side) is replaced by the attention of the Spectraformer which pre-computes the Hadamard (element-wise) product of values $\bm{v}$ and transformed keys $\phi(\bm{k})$ scaled by $\phi(\bm{k})$, the product of which is multiplied with the transformed queries $\phi(\bm{q})$. The linear map $\phi$ is composed from weight matrices and component functions. The linear map has $s$ components with $a$ being the component weight (see Equation 11). The component weight $a$ can either be fixed (typically at $a=1$) or learnable as proposed by various kernel works (Liu et al., 2022). Due to the fact that the majority of random feature Transformers (e.g., Likhosherstov et al. (2022)) have $a=1$, we follow this setting for consistent comparison between combinations.
The figure and the equation together highlight the three strands of research as mentioned previously in Section 1: (A) Weight matrix approximator: constructing $\bm{W}$ more effectively; (B) Weight matrix learner: parameterizing $\bm{W}$ instead of sampling; (C) Component function: constructing $f_{j}$ more effectively.
@@ 3.1 Weight Matrices
In Spectraformer, the weight matrix $\bm{W}$ which is either an approximator (from the families of (e.g., ORF and SORF (Yu et al., 2016)), Unit-cube (e.g., QMC (Avron et al., 2016), MM (Shen et al., 2017)), Quadrature (e.g., SGQ (Dao et al., 2017)). or a learner (FastFoodL (Chowdhury et al., 2022)).
@@
We term the matrix $\bm{W}$ as weight matrix instead of random matrix since although $\bm{W}$ acts like a random matrix ($\bm{\omega}_{i}\sim p(.)$, see Equation 10), it is not guaranteed to be one. In the case of weight matrix approximator, they approximate a random matrix and in the case of weight matrix learner, they parameterize a weight matrix, thereby implicitly learning a distribution $p(.)$ associated with such $\bm{W}$. The weight matrix is given in Definition 3.1.
@@
Provided that a matrix can substitute for a random matrix in Equation 11, and the solution for this equation given such matrix and $f$ as TrigRF being the solution for Equation 9, i.e., as an unbiased estimator of $\mathcal{K}$, and TrigRF being $f_{1}=f_{l}=\exp(-i\bm{\omega}^{T}\bm{x})$ (see Equation 10) then such a matrix is a weight matrix.
@@
It follows from Definition 3.1 that for any weight matrix $\bm{W}$, $f=TrigRF$ is a valid component function. Spectraformer enables weight matrix approximation in terms of three families based on how they solve the intractable integral in Equation 9:
@@
  • Monte Carlo sampling-based weight matrix involves approximating the integral of Equation 9 using Monte Carlo sampling with the solution given by Equation 10 (Rahimi & Recht, 2007) with $\bm{W}$ being the ‘Base’ random matrix of $p(.)$, $\bm{\omega}_{i}\sim p(.)$. The random matrix $\bm{W}$ can be constructed more efficiently (either reducing time or space complexity) with FastFood (Le et al., 2013). The random matrix $\bm{W}$ can also be enforced to form geometrical couplings such as orthogonality to reduce approximation error with ORF, SORF (Yu et al., 2016). Both these methods produce valid random matrices.
@@
  • Unit-cube sampling-based weight matrix transforms Equation 9 to an integral on the unit cube $[0,1]^{d}$, then performs the approximation with uniform and independent points, thus reducing variance. The first approach is QMC (Avron et al., 2016), where $\bm{W}:\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$, $\Phi$ being the cumulative distribution function (CDF) associated with $p(.)$ and $\bm{t}_{i}$ being a low discrepancy sequence. $\Phi^{-1}(\bm{t}_{i})\sim p(.)$ whilst having a lower variance than direct sampling from $p(.)$. MM (Shen et al., 2017; Liu et al., 2022) improves over QMC by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$ on $\Phi^{-1}$.
@@
  • Quadrature-based weight matrix uses quadrature rules with non-uniform deterministic weights. The main approach explored is SGQ (Dao et al., 2017) which uses one-dimensional Gaussian quadrature rule by assuming $\mathcal{K}$ factorizes with respect to the dimensions. Smolyak rule is further implemented to alleviate the curse of dimensionality.
@@
Liu et al. (2022) show that in various kernel benchmark datasets (among combination with TrigRF), QMC and ORF have consistently high performance. Due to the significant number of weight matrices in random features literature, we decide to cover only the most prominent ones. Appendix A.3 provides the technical details of the approximators discussed here and those that we left out.
@@
Spectraformer also allows to learn weight matrices by learning $p(.)$ via parameterizing $\bm{W}$. Yang et al. (2015) introduce several of these parametrization schemes, which are then adapted into the attention setting by Chowdhury et al. (2022). The best performing method is via making the constituent matrices of FastFood learnable, which we term FastFoodL. In general, a weight matrix learner is valid as long as it can implicitly model any distribution. In addition the TrigRF, PosRF has also been applied on FastFoodL. Technical details and other learners we have not covered are explored in Appendix A.4. We now explore existing works which have developed from Rahimi & Recht (2007) (see Equation 9, 10) and how they lead to the development of Spectraformer.
@@ 3.2 Component functions
Spectraformer incorporates component functions $f$ such as PosRF, OPRF, SADERF. Component functions in Spectraformer combine each weight matrix row $\bm{\omega}_{i}$ with the input $\bm{x}$. The base case of the component function is TrigRF. Component function is given in Definition 3.2
@@
A function of $\bm{x}$ with respect to $\bm{\omega}$ is a valid component function if and only if $\mathcal{K}\approx\phi_{\bm{\omega}}(\bm{x})^{T}\phi_{\bm{\omega}}(\bm{y})$ given that $\bm{\omega}_{i}\sim p(.)$ in Equation 10.
@@
It follows that all component functions have provided theoretical guarantees given that $\bm{W}$ is a random matrix. In addition, PosRF, OPRF, and SADERF have also been proven theoretically to approximate $\mathcal{K}_{RBF}$ given $\bm{W}$ being the ORF.
@@
While TrigRF(Rahimi & Recht, 2007) is the base component function and guarantees $\mathcal{K}(\bm{\delta})$ to be close to $\phi_{TrigRF}(\bm{x})\cdot\phi_{TrigRF}(\bm{y})$ due to uniform convergence, it is typically evaluated using only the real component of $\exp(-i\bm{\omega}_{j}^{T}\bm{x})$. This leads to $\cos$ and $\sin$ functions which can produce unstable behavior, especially when $K(\bm{\delta})$ is close to 0, where variance is infinite (Choromanski et al., 2021; Likhosherstov et al., 2022). Therefore, TrigRF is avoided in Spectraformer.
@@
  • PosRF (Choromanski et al., 2021) is introduced in Spectraformer as a simple modification over TrigRF to counter this instability. PosRF guarantees positive function output and its variance approaching 0 as $K(\bm{\delta})$ approaches 0. Likhosherstov et al. (2022) generalize TrigRF and PosRF to GERF.
@@
  • OPRF (Likhosherstov et al., 2022) is one of the solution to the reduction of the variance equation of GERF. OPRF is both positive and bounded (whereas PosRF is positive and unbounded). Likhosherstov et al. (2023) generalize GERF to dense matrix instead of scalar parameters leading to DERF.
@@
  • SADERF (Likhosherstov et al., 2023) is one of the solutions to the reduction of the variance equation of DERF. Whilst DERF has many solutions, they all rely on operations which are incompatible with many deep learning libraries on GPU and TPU like SVD and eigen decompositions. SADERF is an extension of GERF with a tighter variance bound on DERF.
@@
The alternatives for component functions allow Spectraformer to approximate the RBF with $p(.)$ corresponding to the Gaussian. Likhosherstov et al. (2023) and Likhosherstov et al. (2022) show that SADERF-ORF is the SOTA combination followed by OPRF-ORF and PosRF-ORF. Technical details and other component function(s) not covered are explored in Appendix A.5.
@@
Table 2: Summary of research gap in kernelized attention. Rows are the component functions, columns are the weight matrices. For each cell, the top item is the cited work for this combination for typical kernel tasks, while the bottom item is the cited work for Transformer-based tasks. $\times$ indicates no prior work for the combination. Spectraformer explores all combinations listed here, excluding ‘Base’ due to exhaustive work as seen in the table.
TrigRF PosRF OPRF SADERF
Monte Carlo Base (Rahimi & Recht, 2007) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
(Choromanski et al., 2021) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
ORF (Yu et al., 2016) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
(Choromanski et al., 2021) (Choromanski et al., 2021) (Likhosherstov et al., 2022) (Likhosherstov et al., 2023)
SORF (Yu et al., 2016) $\times$ $\times$ $\times$
$\times$ $\times$ $\times$ $\times$
Unit-cube QMC (Avron et al., 2016) $\times$ $\times$ $\times$
$\times$ $\times$ $\times$ $\times$
MM (Shen et al., 2017) $\times$ $\times$ $\times$
$\times$ $\times$ $\times$ $\times$
Quadrature SGQ (Dao et al., 2017) $\times$ $\times$ $\times$
$\times$ $\times$ $\times$ $\times$
Learner FastFoodL (Yang et al., 2015) $\times$ $\times$ $\times$
(Chowdhury et al., 2022) (Chowdhury et al., 2022) $\times$ $\times$